Broward Manufacturing recently reported the following information: Net income $765,000 ROA 10% Interest expense $229,500 Accounts payable and accruals $1,050,000 Broward's tax rate is 25%. Broward finances with only debt and common equity, so it has no preferred stock. 40% of its total invested capital is debt, and 60% of its total invested capital is common equity. Calculate its basic earning power (BEP), its return on equity (ROE), and its return on invested capital (ROIC). Do not round intermediate calculations. Round your answers to two decimal places.

BEP:

ROE:

ROIC:
Answer #1

Answer:

BEP = 16.33%

ROE = 19.32%

ROIC = 14.20%

Working:

BEP:

Given:

Net Income = $765,000

Interest expense = $229,500

Broward's tax rate is = 25%

ROA = 10%

Income before tax =  Net Income / (1 - Tax rate) = 765000/ (1 - 25%) = 765000/ 75% = $1,020,000

EBIT = Income before tax + Interest expense = 1020000 + 229500 = $1,249,500

Total assets = Net Income / ROA = 765000 / 10% = $7,650,000

Basic Earning Power = EBIT / Total assets = $1,249,500 / 7,650,000 = 16.33%

ROE:

Invested capital = Total assets - Accounts payable and accruals = 7650000 - 1050000 = $6,600,000

60% of its total invested capital is common equity.

Equity = Invested capital * 60% = 6600000 * 60% = $3.960,000

ROE = Net Income / Equity = 765000 / 3,960,000 = 19.32%

ROIC:

ROIC = NOPAT / Invested capital = EBIT * (1 - tax rate) / Invested capital

= 1,249,500 * (1 - 25%) / 6600000

= 14.20%